Jefferson Lab Weekly Briefs
June 13, 2012
12 GeV Upgrade
The full research performance goal of the C100 cryomodules has been successfully demonstrated: 108 Megavolts while accelerating 465 µA (microamps) for more than one hour. The major installation work is well underway. The east recombiner has been removed and the removal of the west spreader has begun. Two new cryomodules (C100-4 and C100-5) have been installed. More than ten trim power supply racks have been installed. Installation of the compressors and 4.5 Kelvin coldbox for the Central Helium Liquefier-2 continues. The beam position monitors for arc 10 are ready for installation.
Staffers from all three halls have been consumed with various meetings, reviews and other items, including last week's JLab Users Group Annual Meeting and Workshop, followed by collaboration meetings for all three halls in June, plus other meetings.
This week was the final full week of running electron beam in the injector to support the Polarized Electrons for Polarized Positrons (PEPPo) experiment. The radiofrequency capture section, which had stability issues at the end of last week, remained stable, but Engineering Support personnel maintained a close eye on it. On Thursday afternoon, an anomaly with a cryogenic valve position indication caused a pressure surge that shut down the CHL, forcing PEPPo to pause for a few hours. Later that day, the South Linac tunnel was locked up in preparation for commissioning of the superconducting RF cryogenic module 2L23. The weekend was fairly quiet running. Early Sunday morning, the PEPPo beamline experienced a sudden change in vacuum. Investigation of the beamline and components revealed no reason for the change. A leak check found no leaks, and an analysis of the gasses inside the beamline showed the gasses weren't from the surrounding air, indicating the source of the drop in vacuum may have been from something outgassing within the beamline. The vacuum recovered and PEPPo was able to continue with little interruption until the accelerator was shut down and completely opened up for access for 12 GeV work on Monday morning.
The new Argonne-undulator-based ultraviolet FEL has been running well, and last week the FEL team optimized the lasing setup to produce 124 nanometer photons in the third harmonic. This photon beam was then brought up into User Lab 1, and its power and spectrum were characterized. On Friday, the user chamber for the Atomic Trap Trace Analysis work on krypton was connected to the vacuum system. In addition, the team continued re-work of the Darklight hardware to prepare it for fiducialization and installation in the coming weeks.
With the start of the 12 GeV installation shutdown about two weeks ago, the Electrical Engineering Systems (EES) group has switched from largely accelerator operations support to multiple shutdown activities. Power supplies and RF systems have been shut down and put in a safe mode so that other tasks can begin. The magnet power supplies that are no longer needed have been removed to storage. The equipment tear out, for upgrades, of many accelerator components has required disconnecting thousands of power and signal cables prior to removal of the equipment. This work must be done carefully to prevent damage and to allow correct hook-up once the new hardware is re-installed. Nine EES personnel have been assigned to help other groups during the initial phase of the shutdown. Five people are working in Hall B on equipment and cable removal and another four are working with the Mechanical Installation group to assist in component removal. The safety systems group is engaged in revamping and modifying the Personnel Safety System to support the upgrades. Finally, a small cadre of people continue to support the remaining physics operation: the PEPPo experiment in the injector.
E-Sign System for Documents
E-Sign has been tested throughout the lab in two phases to include novices as well as seasoned users of the current paper system. To ensure a smooth transition from the current system to E-Sign, several training sessions are being scheduled between June 15 and June 30. Training documentation is also available on the M:drive : M:dcgE-Sign Testing. If you have questions about E-Sign, contact Butch Dillon-Townes, x7563.
Facilities Management and Logistics
JLab Energy Reduction Annual Test Set for Thursday, June 14
Environment, Safety, Health and Quality
Your work documents should list specifically what type of glove should be used for the job. If it does not, then it is time to stop and ask your supervisor. If no one knows, then it is time to ask Industrial Hygiene (Jen Williams, x7882) for help. Most gloves are available in the Stock Room for the asking, but make sure you are asking for the correct one.
Computing and Networking Infrastructure
CNI Maintenance Period Set for June 19, 5-7 p.m.
Email Migrations into New System Continue
New Montage on the Month of May
Education Seeks Projects for Summer Honors Interns
Tracking Thomas at JLab
Congratulations this week go to Samika Hawkins, who was the only person to correctly identify the June 6 location. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.
JLab Calendar of Events
June 14: Accelerator Colloquium and Public Lecture: ADS and the MYRRHA project